OpenStudy (anonymous):

MEDAL AND FAN!! The elevation of city A is 60 feet above sea level; the elevation of city C is 12 feet above sea level. The elevation of city D is 1/2 of the elevation of city C. Part A: Write an expression to represent the elevation, in feet, of city D. Part B: Solve the expression and explain why the answer is negative or positive. @Princer_Jones

OpenStudy (anonymous):

Whats part A?

OpenStudy (anonymous):

I think a phrase is missing in the question. Write the last sentence again...

OpenStudy (studygurl14):

There seems to be a missing part of the sentence "The elevation of city D is ___ of the elevation of city C"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I did it

OpenStudy (anonymous):

@StudyGurl14 you help her. It is very easy.:)

OpenStudy (anonymous):

why cant you help me? your give me great answers?

OpenStudy (anonymous):

can you please help me? @Princer_Jones

OpenStudy (anonymous):

Ok. Elevation of city D is half the elevation of city C, It is given that the evelation of city C is 12 feet above the sea level level. Hence elevation of city D=12/2=6feet above the sea level

OpenStudy (anonymous):

So whats the expression for Part A?

OpenStudy (anonymous):

The elevation of city D=x/2, where x is the elevation of city C in feet

OpenStudy (anonymous):

For part B write, here x=12 hence x/2=6 Therefore the elevation of city D is 6feet above the sea level

OpenStudy (anonymous):

And is the answer negative or postive?

OpenStudy (anonymous):

positive since when a positive number is divided by a positive number, the solution is positive.

OpenStudy (anonymous):

ok! That was awesome! I cant thank you enough :) Can u help me with some more pretty please? Ill keep giving you medals :) so yea?

OpenStudy (anonymous):

can you help me with more please?? :) @Princer_Jones

OpenStudy (anonymous):

you tag me in. I am more ten minutes, .. so do it quick, okay?:)

OpenStudy (anonymous):

ok :)